\begin{thebibliography}{999}

\bibitem{Cornwall:1981zr}  J.~M.~Cornwall,
%``Dynamical Mass Generation In Continuum QCD,''
Phys.\ Rev.\ D \textbf{26}, 1453 (1982).

%\cite{Greensite:1985vq}
\bibitem{Greensite:1985vq}  J.~Greensite and M.~B.~Halpern,
%``Variational Computation Of Glueball Masses In Continuum QCD,''
Nucl.\ Phys.\ B \textbf{271}, 379 (1986).

%\cite{Stingl:1985hx}

\bibitem{Stingl:1985hx}  M.~Stingl,
%``Propagation Properties And Condensate Formation Of The Confined Yang-Mills
%Field,''
Phys.\ Rev.\ D \textbf{34}, 3863 (1986) [Erratum-ibid.\ D
\textbf{36}, 651 (1987)].

%\cite{Lavelle:1988eg}

\bibitem{Lavelle:1988eg}  M.~J.~Lavelle and M.~Schaden,
%``Propagators And Condensates In QCD,''
Phys.\ Lett.\ B \textbf{208}, 297 (1988).

%\cite{Gubarev:2000nz}

\bibitem{Gubarev:2000nz}  F.~V.~Gubarev and V.~I.~Zakharov,
%``On the emerging phenomenology of <(A(a)(mu))**2(min)>,''
Phys.\ Lett.\ B \textbf{501}, 28 (2001) [arXiv:hep-ph/0010096].

%\cite{Gubarev:2000eu}

\bibitem{Gubarev:2000eu}  F.~V.~Gubarev, L.~Stodolsky and V.~I.~Zakharov,
%``On the significance of the quantity A**2,''
Phys.\ Rev.\ Lett.\ \textbf{86}, 2220 (2001)
[arXiv:hep-ph/0010057].
%%CITATION = HEP-PH 0010057;%%

%\bibitem{Verschelde:2001ia,Dudal:2003vv,Browne:2003uv}

\bibitem{Verschelde:2001ia}  H.~Verschelde, K.~Knecht, K.~Van Acoleyen and
M.~Vanderkelen,
%``The non-perturbative groundstate of QCD and the local composite  operator
%A(mu)**2,''
Phys.\ Lett.\ B \textbf{516}, 307 (2001) [arXiv:hep-th/0105018].

%\cite{Kondo:2001nq}

\bibitem{Kondo:2001nq}  K.~I.~Kondo,
%``Vacuum condensate of mass dimension 2 as the origin of mass gap and  quark
%confinement,''
Phys.\ Lett.\ B \textbf{514}, 335 (2001) [arXiv:hep-th/0105299].
%%CITATION = HEP-TH 0105299;%%

%\cite{Kondo:2001tm}

\bibitem{Kondo:2001tm}  K.~I.~Kondo, T.~Murakami, T.~Shinohara and T.~Imai,
%``Renormalizing a BRST-invariant composite operator of mass dimension 2  in
%Yang-Mills theory,''
Phys.\ Rev.\ D \textbf{65}, 085034 (2002) [arXiv:hep-th/0111256].
%%CITATION = HEP-TH 0111256;%%

%\cite{Dudal:2003vv}

\bibitem{Dudal:2003vv}  D.~Dudal, H.~Verschelde, R.~E.~Browne and
J.~A.~Gracey,
%``A determination of  and the non-perturbative vacuum energy of
%Yang-Mills theory in the Landau gauge,''
Phys.\ Lett.\ B \textbf{562}, 87 (2003) [arXiv:hep-th/0302128].
%%CITATION = HEP-TH 0302128;%%

%\cite{Browne:2003uv}

\bibitem{Browne:2003uv}  R.~E.~Browne and J.~A.~Gracey,
%``Two loop effective potential for  in the Landau gauge in  quantum
%chromodynamics,''
JHEP \textbf{0311}, 029 (2003) [arXiv:hep-th/0306200].
%%CITATION = HEP-TH 0306200;%%

%\cite{Dudal:2003gu}

\bibitem{Dudal:2003gu}  D.~Dudal, H.~Verschelde, V.~E.~R.~Lemes,
M.~S.~Sarandy, S.~P.~Sorella and M.~Picariello,
%``Gluon-ghost condensate of mass dimension 2 in the Curci-Ferrari gauge,''
Annals Phys.\ \textbf{308}, 62 (2003) [arXiv:hep-th/0302168].
%%CITATION = HEP-TH 0302168;%%

%\cite{Dudal:2003by}

\bibitem{Dudal:2003by}  D.~Dudal, H.~Verschelde, J.~A.~Gracey,
V.~E.~R.~Lemes, M.~S.~Sarandy, R.~F.~Sobreiro and S.~P.~Sorella,
%``Dynamical gluon mass generation from  in linear covariant gauges,''
JHEP \textbf{0401}, 044 (2004) [arXiv:hep-th/0311194].
%%CITATION = HEP-TH 0311194;%%

%\cite{Dudal:2004rx}

\bibitem{Dudal:2004rx}  D.~Dudal, J.~A.~Gracey, V.~E.~R.~Lemes,
M.~S.~Sarandy, R.~F.~Sobreiro, S.~P.~Sorella and H.~Verschelde,
%``An analytic study of the off-diagonal mass generation for Yang-Mills theories
%in the maximal Abelian gauge,''
Phys.\ Rev.\ D \textbf{70}, 114038 (2004) [arXiv:hep-th/0406132].

%\cite{Browne:2004mk}

\bibitem{Browne:2004mk}  R.~E.~Browne and J.~A.~Gracey,
%``One loop MS-bar gluon pole mass from the LCO formalism,''
Phys.\ Lett.\ B \textbf{597}, 368 (2004) [arXiv:hep-ph/0407238].
%%CITATION = HEP-PH 0407238;%%

%\cite{Gracey:2004bk}

\bibitem{Gracey:2004bk}  J.~A.~Gracey,
%``Two loop MS-bar gluon pole mass from the LCO formalism,''
Eur.\ Phys.\ J.\ C \textbf{39}, 61 (2005) [arXiv:hep-ph/0411169].
%%CITATION = HEP-PH 0411169;%%

%\cite{Li:2004te}

\bibitem{Li:2004te}  X.~d.~Li and C.~M.~Shakin,
%``Description of gluon propagation in the presence of an A**2 condensate,''
Phys.\ Rev.\ D \textbf{71}, 074007 (2005) [arXiv:hep-ph/0410404].
%%CITATION = HEP-PH 0410404;%%

%\cite{Boucaud:2001st}

\bibitem{Boucaud:2001st}  P.~Boucaud, A.~Le Yaouanc, J.~P.~Leroy,
J.~Micheli, O.~Pene and J.~Rodriguez-Quintero,
%``Testing Landau gauge OPE on the lattice with a  condensate,''
Phys.\ Rev.\ D \textbf{63}, 114003 (2001) [arXiv:hep-ph/0101302].
%%CITATION = HEP-PH 0101302;%%

%\cite{Boucaud:2002nc}

\bibitem{Boucaud:2002nc}  P.~Boucaud, J.~P.~Leroy, A.~Le Yaouanc, J.~Micheli, O.~Pene, F.~De Soto, A.~Donini, H.~Moutare and J.~Rodriguez-Quintero
%``Instantons and  condensate,''
Phys.\ Rev.\ D \textbf{66}, 034504 (2002) [arXiv:hep-ph/0203119].
%%CITATION = HEP-PH 0203119;%%

%\cite{Boucaud:2005rm}

\bibitem{Boucaud:2005rm}  P.~Boucaud, F.~de Soto, J.~P.~Leroy, A.~Le Yaouanc, J.~Micheli, H.~Moutarde,
O.~Pene and J.~Rodriguez-Quintero, %``Artefacts and  power corrections: Revisiting the MOM Z(psi)(p**2) and
%Z(V),''
arXiv:hep-lat/0504017. %%CITATION = HEP-LAT 0504017;%%

%\cite{RuizArriola:2004en}

\bibitem{RuizArriola:2004en}  E.~Ruiz Arriola, P.~O.~Bowman and
W.~Broniowski,
%``Landau-gauge condensates from the quark propagator on the lattice,''
Phys.\ Rev.\ D \textbf{70}, 097505 (2004) [arXiv:hep-ph/0408309].
%%CITATION = HEP-PH 0408309;%%

%\cite{Suzuki:2004dw}

\bibitem{Suzuki:2004dw}  T.~Suzuki, K.~Ishiguro, Y.~Mori and T.~Sekido,
%``The dual Meissner effect and Abelian magnetic displacement currents,''
Phys.\ Rev.\ Lett.\ \textbf{94}, 132001 (2005)
[arXiv:hep-lat/0410001].
%%CITATION = HEP-LAT 0410001;%%

%\cite{Gubarev:2005it}

\bibitem{Gubarev:2005it}  F.~V.~Gubarev and S.~M.~Morozov,
%`` condensate, Bianchi identities and chromomagnetic fields  degeneracy
%in SU(2) YM theory,''
Phys.\ Rev.\ D \textbf{71}, 114514 (2005) [arXiv:hep-lat/0503023].
%%CITATION = HEP-LAT 0503023;%%

%\cite{Furui:2005bu}


\bibitem{Furui:2005bu}
S. Furui and H. Nakajima, Hideo,
%"Infrared features of unquenched Lattice Landau Gauge QCD",
Few Body Syst. 40, 101-128 (2006) [arXiv:hep-lat/0503029].



%\cite{Boucaud:2005xn}

\bibitem{Boucaud:2005xn}  P.~Boucaud, J.~P.~Leroy, A.~Le Yaouanc, A.~Y.~Lokhov, J.~Micheli, O.~Pene,
J.~Rodriguez-Quintero and C.~Roiesnel,
%``Non-perturbative power corrections to ghost and gluon propagators,''
arXiv:hep-lat/0507005. %%CITATION = HEP-LAT 0507005;%%

%\cite{Chernodub:2005gz}

\bibitem{Chernodub:2005gz}  M.~N.~Chernodub, K.~Ishiguro, Y.~Mori, Y.~Nakamura, M.~I.~Polikarpov,
T.~Sekido, T.~Suzuki and V.~I.~Zakharov,
%``Vacuum type of SU(2) gluodynamics in maximally Abelian and Landau gauges,''
arXiv:hep-lat/0508004. %%CITATION = HEP-LAT 0508004;%%

%\cite{Boucaud:2005xn,Boucaud:2008gn,Pene:2011kg,Boucaud:2010gr,Blossier:2010ky,Boucaud:2011eh,Blossier:2011tf,Blossier:2013te}

%\cite{Boucaud:2005xn}
\bibitem{Boucaud:2005xn}
  P.~Boucaud, J.~P.~Leroy, A.~Le Yaouanc, A.~Y.~Lokhov, J.~Micheli, O.~Pene, J.~Rodriguez-Quintero and C.~Roiesnel,
  %``Non-perturbative power corrections to ghost and gluon propagators,''
  JHEP {\bf 0601}, 037 (2006)
  doi:10.1088/1126-6708/2006/01/037
  [hep-lat/0507005].
  %%CITATION = doi:10.1088/1126-6708/2006/01/037;%%
  %36 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Boucaud:2008gn}
\bibitem{Boucaud:2008gn}
  P.~Boucaud, F.~De Soto, J.~P.~Leroy, A.~Le Yaouanc, J.~Micheli, O.~Pene and J.~Rodriguez-Quintero,
  %``Ghost-gluon running coupling, power corrections and the determination of Lambda(MS-bar),''
  Phys.\ Rev.\ D {\bf 79}, 014508 (2009)
  doi:10.1103/PhysRevD.79.014508
  [arXiv:0811.2059 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.79.014508;%%
  %70 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Pene:2011kg}
\bibitem{Pene:2011kg}
  O.~Pene {\it et al.},
  %``Vacuum expectation value of $A^2$ from LQCD,''
  PoS FACESQCD {\bf }, 010 (2010)
  [arXiv:1102.1535 [hep-lat]].
  %%CITATION = ARXIV:1102.1535;%%
  %5 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Boucaud:2010gr}
\bibitem{Boucaud:2010gr}
  P.~Boucaud, M.~E.~Gomez, J.~P.~Leroy, A.~Le Yaouanc, J.~Micheli, O.~Pene and J.~Rodriguez-Quintero,
  %``The low-momentum ghost dressing function and the gluon mass,''
  Phys.\ Rev.\ D {\bf 82}, 054007 (2010)
  doi:10.1103/PhysRevD.82.054007
  [arXiv:1004.4135 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.82.054007;%%
  %38 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Blossier:2010ky}
\bibitem{Blossier:2010ky}
  B.~Blossier {\it et al.} [ETM Collaboration],
  %``Ghost-gluon coupling, power corrections and $\Lambda_{\overline {\rm MS}}$ from twisted-mass lattice QCD at Nf=2,''
  Phys.\ Rev.\ D {\bf 82}, 034510 (2010)
  doi:10.1103/PhysRevD.82.034510
  [arXiv:1005.5290 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.82.034510;%%
  %41 citations counted in INSPIRE as of 19 Apr 2016
  
  %\cite{Dudal:2010tf}
\bibitem{Dudal:2010tf} 
  D.~Dudal, O.~Oliveira and N.~Vandersickel,
  %``Indirect lattice evidence for the Refined Gribov-Zwanziger formalism and the gluon condensate $\langle{A^2}\rangle$ in the Landau gauge,''
  Phys.\ Rev.\ D {\bf 81}, 074505 (2010)
  doi:10.1103/PhysRevD.81.074505
  [arXiv:1002.2374 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.81.074505;%%
  %113 citations counted in INSPIRE as of 30 Jul 2016

  %\cite{Boucaud:2011eh}
\bibitem{Boucaud:2011eh}
  P.~Boucaud, D.~Dudal, J.~P.~Leroy, O.~Pene and J.~Rodriguez-Quintero,
  %``On the leading OPE corrections to the ghost-gluon vertex and the Taylor theorem,''
  JHEP {\bf 1112}, 018 (2011)
  doi:10.1007/JHEP12(2011)018
  [arXiv:1109.3803 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP12(2011)018;%%
  %19 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Blossier:2011tf}
\bibitem{Blossier:2011tf}
  B.~Blossier {\it et al.},
  %``Ghost-gluon coupling, power corrections and $\Lambda_{\bar{\rm MS}}$ from lattice QCD with a dynamical charm,''
  Phys.\ Rev.\ D {\bf 85}, 034503 (2012)
  doi:10.1103/PhysRevD.85.034503
  [arXiv:1110.5829 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.85.034503;%%
  %28 citations counted in INSPIRE as of 19 Apr 2016

  %\cite{Blossier:2013te}
\bibitem{Blossier:2013te}
  B.~Blossier, P.~Boucaud, M.~Brinet, F.~De Soto, V.~Morenas, O.~Pene, K.~Petrov and J.~Rodriguez-Quintero,
  %``Testing the OPE Wilson coefficient for $A^2$ from lattice QCD with a dynamical charm,''
  Phys.\ Rev.\ D {\bf 87}, 074033 (2013)
  doi:10.1103/PhysRevD.87.074033
  [arXiv:1301.7593 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.87.074033;%%
  %9 citations counted in INSPIRE as of 19 Apr 2016



\bibitem{Dudal:2002pq}  D.~Dudal, H.~Verschelde and S.~P.~Sorella,
%``The anomalous dimension of the composite operator A**2 in the Landau
%gauge,''
Phys.\ Lett.\ B \textbf{555}, 126 (2003) [arXiv:hep-th/0212182].
%%CITATION = HEP-TH 0212182;%%


%\cite{Gracey:2002yt}
\bibitem{Gracey:2002yt}
  J.~A.~Gracey,
  %``Three loop MS-bar renormalization of the Curci-Ferrari model and the
  %dimension two BRST invariant composite operator in QCD,''
  Phys.\ Lett.\ B {\bf 552} (2003) 101
  [arXiv:hep-th/0211144].
  %%CITATION = HEP-TH 0211144;%%



\bibitem{Dudal:2005na}  D.~Dudal, R.~F.~Sobreiro, S.~P.~Sorella and
H.~Verschelde,
%``The Gribov parameter and the dimension two gluon condensate in Euclidean
%Yang-Mills theories in the Landau gauge,''
Phys.\ Rev.\ D \textbf{72}, 014016 (2005) [arXiv:hep-th/0502183].
%%CITATION = HEP-TH 0502183;%%

%\cite{Dudal:2007cw}
\bibitem{Dudal:2007cw}
  D.~Dudal, S.~P.~Sorella, N.~Vandersickel and H.~Verschelde,
  %``New features of the gluon and ghost propagator in the infrared region from the Gribov-Zwanziger approach,''
  Phys.\ Rev.\ D {\bf 77}, 071501 (2008)
  doi:10.1103/PhysRevD.77.071501
  [arXiv:0711.4496 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.77.071501;%%
  %157 citations counted in INSPIRE as of 28 Apr 2016

%\cite{Dudal:2008sp}
\bibitem{Dudal:2008sp}
  D.~Dudal, J.~A.~Gracey, S.~P.~Sorella, N.~Vandersickel and H.~Verschelde,
  %``A Refinement of the Gribov-Zwanziger approach in the Landau gauge: Infrared propagators in harmony with the lattice results,''
  Phys.\ Rev.\ D {\bf 78} (2008) 065047
  doi:10.1103/PhysRevD.78.065047
  [arXiv:0806.4348 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.78.065047;%%
  %313 citations counted in INSPIRE as of 28 Apr 2016

%\cite{Dudal:2011gd}
\bibitem{Dudal:2011gd}
  D.~Dudal, S.~P.~Sorella and N.~Vandersickel,
  %``The dynamical origin of the refinement of the Gribov-Zwanziger theory,''
  Phys.\ Rev.\ D {\bf 84}, 065039 (2011)
  doi:10.1103/PhysRevD.84.065039
  [arXiv:1105.3371 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.84.065039;%%
  %70 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Vandersickel:2012tz}
\bibitem{Vandersickel:2012tz}
  N.~Vandersickel and D.~Zwanziger,
  %``The Gribov problem and QCD dynamics,''
  Phys.\ Rept.\  {\bf 520} (2012) 175
  doi:10.1016/j.physrep.2012.07.003
  [arXiv:1202.1491 [hep-th]].
  %%CITATION = doi:10.1016/j.physrep.2012.07.003;%%
  %101 citations counted in INSPIRE as of 28 Apr 2016


 % \cite{Aguilar:2008xm,Aguilar:2015bud}

  %\cite{Aguilar:2008xm}
\bibitem{Aguilar:2008xm}
  A.~C.~Aguilar, D.~Binosi and J.~Papavassiliou,
  %``Gluon and ghost propagators in the Landau gauge: Deriving lattice results from Schwinger-Dyson equations,''
  Phys.\ Rev.\ D {\bf 78} (2008) 025010
  doi:10.1103/PhysRevD.78.025010
  [arXiv:0802.1870 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.78.025010;%%
  %333 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Aguilar:2015bud}
\bibitem{Aguilar:2015bud}
  A.~C.~Aguilar, D.~Binosi and J.~Papavassiliou,
  %``The Gluon Mass Generation Mechanism: A Concise Primer,''
  Front.\ Phys.\ China {\bf 11} (2016) no.2,  111203
  doi:10.1007/s11467-015-0517-6
  [arXiv:1511.08361 [hep-ph]].
  %%CITATION = doi:10.1007/s11467-015-0517-6;%%
  %3 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Fischer:2008uz}
\bibitem{Fischer:2008uz}
  C.~S.~Fischer, A.~Maas and J.~M.~Pawlowski,
  %``On the infrared behavior of Landau gauge Yang-Mills theory,''
  Annals Phys.\  {\bf 324} (2009) 2408
  doi:10.1016/j.aop.2009.07.009
  [arXiv:0810.1987 [hep-ph]].
  %%CITATION = doi:10.1016/j.aop.2009.07.009;%%
  %318 citations counted in INSPIRE as of 28 Apr 2016




  %\cite{Cucchieri:2007md}
\bibitem{Cucchieri:2007md}
  A.~Cucchieri and T.~Mendes,
  %``What's up with IR gluon and ghost propagators in Landau gauge? A puzzling answer from huge lattices,''
  PoS LAT {\bf 2007}, 297 (2007)
  [arXiv:0710.0412 [hep-lat]].
  %%CITATION = ARXIV:0710.0412;%%
  %269 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Cucchieri:2007rg}
\bibitem{Cucchieri:2007rg}
  A.~Cucchieri and T.~Mendes,
  %``Constraints on the IR behavior of the gluon propagator in Yang-Mills theories,''
  Phys.\ Rev.\ Lett.\  {\bf 100}, 241601 (2008)
  doi:10.1103/PhysRevLett.100.241601
  [arXiv:0712.3517 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevLett.100.241601;%%
  %248 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Cucchieri:2011ig}
\bibitem{Cucchieri:2011ig}
  A.~Cucchieri, D.~Dudal, T.~Mendes and N.~Vandersickel,
  %``Modeling the Gluon Propagator in Landau Gauge: Lattice Estimates of Pole Masses and Dimension-Two Condensates,''
  Phys.\ Rev.\ D {\bf 85} (2012) 094513
  doi:10.1103/PhysRevD.85.094513
  [arXiv:1111.2327 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.85.094513;%%
  %87 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Oliveira:2012eh}
\bibitem{Oliveira:2012eh}
  O.~Oliveira and P.~J.~Silva,
  %``The lattice Landau gauge gluon propagator: lattice spacing and volume dependence,''
  Phys.\ Rev.\ D {\bf 86}, 114513 (2012)
  doi:10.1103/PhysRevD.86.114513
  [arXiv:1207.3029 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.86.114513;%%
  %56 citations counted in INSPIRE as of 28 Apr 2016





\bibitem{Sobreiro:2005vn}  R.~F.~Sobreiro and S.~P.~Sorella,
%``A study of the Gribov copies in linear covariant gauges in Euclidean
%Yang-Mills theories,''
JHEP \textbf{0506}, 054 (2005) [arXiv:hep-th/0506165].
%%CITATION = HEP-TH 0506165;%%


  %\cite{Aguilar:2015nqa}
\bibitem{Aguilar:2015nqa}
  A.~C.~Aguilar, D.~Binosi and J.~Papavassiliou,
  %``Yang-Mills two-point functions in linear covariant gauges,''
  Phys.\ Rev.\ D {\bf 91}, no. 8, 085014 (2015)
  doi:10.1103/PhysRevD.91.085014
  [arXiv:1501.07150 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.91.085014;%%
  %19 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Huber:2015ria}
\bibitem{Huber:2015ria}
  M.~Q.~Huber,
  %``Gluon and ghost propagators in linear covariant gauges,''
  Phys.\ Rev.\ D {\bf 91}, no. 8, 085018 (2015)
  doi:10.1103/PhysRevD.91.085018
  [arXiv:1502.04057 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.91.085018;%%
  %12 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Capri:2015pja}
\bibitem{Capri:2015pja}
  M.~A.~L.~Capri, A.~D.~Pereira, R.~F.~Sobreiro and S.~P.~Sorella,
  %``Non-perturbative treatment of the linear covariant gauges by taking into account the Gribov copies,''
  Eur.\ Phys.\ J.\ C {\bf 75}, no. 10, 479 (2015)
  doi:10.1140/epjc/s10052-015-3707-z
  [arXiv:1505.05467 [hep-th]].
  %%CITATION = doi:10.1140/epjc/s10052-015-3707-z;%%
  %5 citations counted in INSPIRE as of 28 Apr 2016

%\cite{Capri:2015ixa,Capri:2015nzw,Fiorentini:2016rwx}

  %\cite{Capri:2015ixa}
\bibitem{Capri:2015ixa}
  M.~A.~L.~Capri {\it et al.},
  %``Exact nilpotent nonperturbative BRST symmetry for the Gribov-Zwanziger action in the linear covariant gauge,''
  Phys.\ Rev.\ D {\bf 92}, no. 4, 045039 (2015)
  doi:10.1103/PhysRevD.92.045039
  [arXiv:1506.06995 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.92.045039;%%
  %10 citations counted in INSPIRE as of 28 Apr 2016
  
  

  %\cite{Capri:2015nzw}
\bibitem{Capri:2015nzw}
  M.~A.~L.~Capri {\it et al.},
  %``More on the nonperturbative Gribov-Zwanziger quantization of linear covariant gauges,''
  Phys.\ Rev.\ D {\bf 93} (2016) no.6,  065019
  doi:10.1103/PhysRevD.93.065019
  [arXiv:1512.05833 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.93.065019;%%
  %1 citations counted in INSPIRE as of 28 Apr 2016

%\cite{Capri:2016aqq}
\bibitem{Capri:2016aqq} 
  M.~A.~L.~Capri {\it et al.},
  %``A local and BRST-invariant Yang-Mills theory within the Gribov horizon,''
  arXiv:1605.02610 [hep-th].
  %%CITATION = ARXIV:1605.02610;%%

  %\cite{Cucchieri:2009kk}
\bibitem{Cucchieri:2009kk}
  A.~Cucchieri, T.~Mendes and E.~M.~S.~Santos,
  %``Covariant gauge on the lattice: A New implementation,''
  Phys.\ Rev.\ Lett.\  {\bf 103}, 141602 (2009)
  doi:10.1103/PhysRevLett.103.141602
  [arXiv:0907.4138 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevLett.103.141602;%%
  %29 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Cucchieri:2011aa}
\bibitem{Cucchieri:2011aa}
  A.~Cucchieri, T.~Mendes, G.~M.~Nakamura and E.~M.~S.~Santos,
  %``Feynman gauge on the lattice: New results and perspectives,''
  AIP Conf.\ Proc.\  {\bf 1354}, 45 (2011)
  doi:10.1063/1.3587584
  [arXiv:1101.5080 [hep-lat]].
  %%CITATION = doi:10.1063/1.3587584;%%
  %14 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Bicudo:2015rma}
\bibitem{Bicudo:2015rma}
  P.~Bicudo, D.~Binosi, N.~Cardoso, O.~Oliveira and P.~J.~Silva,
  %``Lattice gluon propagator in renormalizable $\xi$ gauges,''
  Phys.\ Rev.\ D {\bf 92}, no. 11, 114514 (2015)
  doi:10.1103/PhysRevD.92.114514
  [arXiv:1505.05897 [hep-lat]].
  %%CITATION = doi:10.1103/PhysRevD.92.114514;%%
  %13 citations counted in INSPIRE as of 28 Apr 2016

  %\cite{Fiorentini:2016rwx}
\bibitem{Fiorentini:2016rwx} 
  M.~A.~L.~Capri, D.~Fiorentini, M.~S.~Guimaraes, B.~W.~Mintz, L.~F.~Palhares and S.~P.~Sorella,
  %``A local and renormalizable framework for the gauge-invariant operator $A^2_{\min}$ in Euclidean Yang-Mills theories in linear covariant gauges,''
  arXiv:1606.06601 [hep-th].
  %%CITATION = ARXIV:1606.06601;%%
  %1 citations counted in INSPIRE as of 30 Jul 2016
  
  \bibitem{Lavelle:1995ty}  M.~Lavelle and D.~McMullan,
%``Constituent quarks from QCD,''
Phys.\ Rept.\ \textbf{279}, 1 (1997) [arXiv:hep-ph/9509344].
%%CITATION = HEP-PH 9509344;%%

   \bibitem{Zwanziger:1990tn}  D.~Zwanziger,
%``Quantization Of Gauge Fields, Classical Gauge Invariance And Gluon
%Confinement,''
Nucl.\ Phys.\ B \textbf{345}, 461 (1990). %%CITATION = NUPHA,B345,461;%%
  
 %\cite{Ferrari:2004pd}
\bibitem{Ferrari:2004pd} 
  R.~Ferrari and A.~Quadri,
  %``Physical unitarity for massive non-Abelian gauge theories in the Landau gauge: Stueckelberg and Higgs,''
  JHEP {\bf 0411}, 019 (2004)
  doi:10.1088/1126-6708/2004/11/019
  [hep-th/0408168].
  %%CITATION = doi:10.1088/1126-6708/2004/11/019;%%
  %24 citations counted in INSPIRE as of 26 May 2016
  
  
  %\cite{Dragon:1996tk}
\bibitem{Dragon:1996tk} 
  N.~Dragon, T.~Hurth and P.~van Nieuwenhuizen,
  %``Polynomial form of the Stuckelberg model,''
  Nucl.\ Phys.\ Proc.\ Suppl.\  {\bf 56B}, 318 (1997)
  doi:10.1016/S0920-5632(97)00341-1
  [hep-th/9703017].
  %%CITATION = doi:10.1016/S0920-5632(97)00341-1;%%
  %13 citations counted in INSPIRE as of 27 May 2016
  
   
 




\end{thebibliography}

\documentclass[3p,times,twocolumn]{elsarticle}
 \biboptions{comma,sort&compress}
 
\usepackage{graphicx}
\usepackage{ecrc}

\volume{00}

\firstpage{1}

\journalname{Nuclear and Particle Physics Proceedings}

\runauth{}

\jid{nppp}

\jnltitlelogo{Nuclear and Particle Physics Proceedings}

\usepackage{amssymb}
\usepackage[figuresright]{rotating}

\begin{document}

\begin{frontmatter}

%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\title{On the  gauge-invariant  operator 
$A^2_{\min}$ in Euclidean Yang-Mills theories}
 % \corref{cor0}}
 \cortext[cor0]{Talk given at 19th International Conference in Quantum Chromodynamics (QCD 16),  4 - 8 july 2016, Montpellier - FR}
 \author[label1]{M.~A.~L.~Capri}
\ead{caprimarcio@gmail.com}
\author[label1]{D.~Fiorentini} 
\ead{diegofiorentinia@gmail.com}
\author[label1]{M.~S.~Guimaraes}
\ead{ msguimaraes@uerj.br}
\author[label1]{B.~W.~Mintz}
\ead{bruno.mintz@uerj.br}    
\author[label1]{L.~F.~Palhares} 
\ead{leticia.palhares@uerj.br} 
\author[label1]{S.~P.~Sorella\fnref{fn1}}
   \fntext[fn1]{Speaker, Corresponding author.}
\ead{silvio.sorella@gmail.com} 
\address[label1]{UERJ -- Universidade do Estado do Rio de Janeiro, Instituto de F\'isica -- Departamento de F\'{\i}sica Te\'orica -- Rua S\~ao Francisco Xavier 524,
20550-013, Maracan\~a, Rio de Janeiro, Brasil}

\pagestyle{myheadings}
\markright{ }
\begin{abstract}
We review our recent work on the gauge-invariant non-local dimension-two operator 
$A^2_{\rm min}$,  whose minimization is defined along the gauge orbit. Albeit non-local, 
the operator $A^2_{\rm min}$ can be  cast in local form through the introduction of an auxiliary Stueckelberg field. 
The whole procedure results into a local action which turns out to be renormalizable to all orders. 
\end{abstract}
% \begin{document}
%\begin{keyword}  
%% keywords here, in the form: keyword \sep keyword

%% MSC codes here, in the form: \MSC code \sep code
%% or \MSC[2008] code \sep code (2000 is the default)

%\end{keyword}

\end{frontmatter}
%%%%%%%%%%%%
%\vspace*{-1.5cm}
\section{Introduction}

The study of operators of dimension two in Yang-Mills theories has already a relatively long history, as confirmed by the 
considerable amount of results obtained through
theoretical and phenomenological studies as well as from lattice
simulations \cite
{Cornwall:1981zr,Greensite:1985vq,Stingl:1985hx,Lavelle:1988eg,Gubarev:2000nz,Gubarev:2000eu,
Verschelde:2001ia,Kondo:2001nq, Kondo:2001tm,Dudal:2003vv,
Browne:2003uv,Dudal:2003gu, Dudal:2003by,
Dudal:2004rx,Browne:2004mk, Gracey:2004bk,
Li:2004te,Boucaud:2001st,Boucaud:2002nc,
Boucaud:2005rm, RuizArriola:2004en,Suzuki:2004dw,Gubarev:2005it,Furui:2005bu,Boucaud:2005xn,
Chernodub:2005gz,Boucaud:2005xn,Boucaud:2008gn,Pene:2011kg,Boucaud:2010gr,Blossier:2010ky,Dudal:2010tf,
Boucaud:2011eh,Blossier:2011tf,Blossier:2013te}.\\\\
In particular, the dimension two gluon condensate $\left\langle A_{\mu }^{a}A_{\mu
}^{a}\right\rangle $ has been much investigated in the Landau gauge. According 
to \cite{Gubarev:2000nz}, this condensate enters the operator
product expansion (OPE) of the gluon propagator. A combined OPE
and lattice analysis has shown that this condensate can account for the $%
1/Q^{2}$ corrections which have been reported \cite
{Boucaud:2001st,Boucaud:2002nc,Boucaud:2005rm,RuizArriola:2004en,Furui:2005bu,
Boucaud:2005xn,Boucaud:2008gn,Pene:2011kg,Boucaud:2010gr,Blossier:2010ky,Boucaud:2011eh,
Blossier:2011tf,Blossier:2013te}
in the running of the coupling constant and in the gluon
correlation functions. \\\\
An effective potential for $\left\langle
A_{\mu }^{a}A_{\mu }^{a}\right\rangle $ in Landau gauge has been obtained and
evaluated in analytic form at two loops in \cite
{Verschelde:2001ia,Dudal:2003vv,Browne:2003uv,Browne:2004mk,Gracey:2004bk},
showing that a nonvanishing value of $\left\langle A_{\mu
}^{a}A_{\mu }^{a}\right\rangle $ is favoured as it lowers the
vacuum energy. As a consequence, a dynamical gluon mass is
generated. We point out that, in the Landau gauge, the operator
$A_{\mu }^{a}A_{\mu }^{a}$ turns out to be $BRST$-invariant on shell, a
property which has allowed for an all-orders proof of its
multiplicative renormalizability  \cite{Dudal:2002pq,Gracey:2002yt}. \\\\
Dimension-two condensates also play an important role within 
the context of the Gribov-Zwanziger approach to confinement 
\cite{Dudal:2005na,Dudal:2007cw,Dudal:2008sp,Dudal:2011gd,Vandersickel:2012tz}  as well as for the 
formation of a dynamical gluon mass within the framework of the Dyson-Schwinger equations in Landau 
gauge, as reported in \cite
{Cornwall:1981zr, Aguilar:2008xm,Aguilar:2015bud}. These non-perturbative effects give rise to the 
so called decoupling solution for the gluon propagator 
\cite{Cornwall:1981zr,Dudal:2005na,Dudal:2007cw,Dudal:2008sp,Aguilar:2008xm,Fischer:2008uz}, 
{\it i.e.} to a propagator which exhibits positivity violation, while attaining a finite non-vanishing 
value at zero momentum. Until now, this behaviour is in very good agreement with the most recent lattice 
numerical simulations  \cite{Cucchieri:2007md,Cucchieri:2007rg,Cucchieri:2011ig,Oliveira:2012eh}. The  
generalization of these results to the linear covariant gauges has been worked out recently and can be found in  
\cite{Sobreiro:2005vn,Aguilar:2015nqa,Huber:2015ria,Capri:2015pja,Capri:2015ixa,Capri:2015nzw,Capri:2016aqq,Cucchieri:2009kk,Cucchieri:2011aa,Bicudo:2015rma}. \\\\Even if a large amount  of results has been obtained, many aspects related to dimension-two operators require further investigation as, for instance,  the issue of the gauge invariance,  a topic of pivotal importance in order to give a precise physical meaning to the corresponding 
condensates. This is precisely the aspect which has been addressed recently in  \cite{Fiorentini:2016rwx} and which will be reviewed in the present contribution.

\section{The gauge invariant operator $A^2_{min}$}

A genuine gauge-invariant dimension-two operator $A_{\min }^{2}$  can be constructed by minimizing  the functional
$\mathrm{Tr}\int d^{4}x\,A_{\mu }^{u}A_{\mu }^{u}$ along the gauge
orbit of $A_{\mu }$, see  \cite{Fiorentini:2016rwx} and refs. therein, namely
\begin{eqnarray}
A_{\min }^{2} &\equiv   &\min_{\{u\}}\mathrm{Tr}{\;}   \frac{1}{2} \int d^{4}x\,A_{\mu
}^{u}A_{\mu }^{u}\;,
\nonumber \\
A_{\mu }^{u} &=&u^{\dagger }A_{\mu }u+\frac{i}{g}u^{\dagger }\partial _{\mu
}u\;.  \label{Aminn0}
\end{eqnarray}
In particular, the stationary condition of the functional (\ref{Aminn0})  gives rise to a non-local transverse 
field configuration $A^h_\mu$, $\partial_\mu A^h_\mu=0$, which can be expressed as an infinite series in 
the gauge field $A_\mu$ \cite{Fiorentini:2016rwx}, {\it i.e.}
\begin{eqnarray}
A_{\mu }^{h} &=& {\cal P}_{\mu \nu} \left( A_{\nu }-ig\left[ \frac{1}{\partial ^{2}}\partial A,A_{\nu
}\right]  \right. \nonumber \\ &+&  \left. \frac{ig}{2}\left[ \frac{1}{\partial ^{2}}\partial A,\partial
_{\nu }\frac{1}{\partial ^{2}}\partial A\right]  \right)  +O(A^{3})      \label{ah}
\end{eqnarray}
where ${\cal P}_{\mu \nu}= (\delta_{\mu\nu} - \frac{\partial_\mu \partial_\nu}{\partial^2})$ is the transverse projector.   
The configuration $A_{\mu }^{h}$ turns out to be left invariant
by infinitesimal gauge transformations order by order in the gauge
coupling $g$ \cite{Lavelle:1995ty} 
\begin{equation}
\delta A_{\mu }^{h} =0\;,  \qquad 
\delta A_{\mu } =-\partial _{\mu }\omega +ig\left[ A_{\mu },\omega \right]
\;.  \label{gio}
\end{equation}
The gauge-invariant nature of expression (\ref{Aminn0})  can be made manifest by rewriting it in terms of the  field
strength $F_{\mu \nu }$. In fact, it turns out that \cite{Zwanziger:1990tn}
\begin{eqnarray}
A_{\min }^{2} = \frac{1}{2} \mathrm{Tr} \int d^4x A^h_\mu A^h_\mu =  -\frac{1}{2}\mathrm{Tr}\int d^{4}x\left( F_{\mu
\nu }\frac{1}{D^{2}}F_{\mu \nu }  \right. \nonumber \\
 + 2i\frac{1}{D^{2}}F_{\lambda \mu
}\left[ \frac{1}{D^{2}}D_{\kappa }F_{\kappa \lambda
},\frac{1}{D^{2}}D_{\nu }F_{\nu \mu }\right]  
\nonumber \\
-2i\left. \frac{1}{D^{2}}F_{\lambda \mu }\left[ \frac{1}{D^{2}}D_{\kappa
}F_{\kappa \nu },\frac{1}{D^{2}}D_{\nu }F_{\lambda \mu }\right] \right)
+O(F^{4})    \nonumber 
\end{eqnarray}
where the operator $({D^{2}})^{-1}$ 
denotes the inverse of the Laplacian $D^2=D_\mu D_\mu$ with $D_\mu$ being the 
covariant derivative \cite{Zwanziger:1990tn}. Let us also notice that, in the particular case of the Landau gauge 
$\partial_\mu A_\mu=0$, the gauge invariant quantity  $(A^h_\mu A^h_\mu)$ reduces to the operator $A^{2}$
\begin{equation}
(A^{h,a}_\mu A^{h,a}_\mu) \Big|_{\rm Landau} =  A^a_\mu A^a_\mu \;.      \label{land}
\end{equation}

\section{Construction of a local Lagrangian for $A^2_{min}$}

In order to construct a local action,  we start 
with the standard Faddeev-Popov action of Yang-Mills theory quantized in 
linear covariant gauges with the inclusion of the non-local gauge invariant  operator $(A^h_\mu A^h_\mu)$ as well as of a constraint enforcing the 
transversality of the field configuration $A_\mu^h$, {\it i.e.} we consider the action 
\begin{equation}
S = S_{FP}+
\int d^4x \left(\tau^{a}\,\partial_{\mu}A^{h,a}_{\mu}
+\frac{m^{2}}{2}\,A^{h,a}_{\mu}A^{h,a}_{\mu}\right)  \label{sact}
\end{equation} 
where $S_{FP}$ stands for the Faddeev-Popov action in linear covariant gauges   
\[
S_{FP} = \int d^{4}x\,\bigg(\frac{1}{4}\,F^{a}_{\mu\nu}F^{a}_{\mu\nu}+\frac{\alpha}{2}\,b^{a}b^{a}
+ib^{a}\,\partial_{\mu}A^{a}_{\mu}
+\bar{c}^{a}\,\partial_{\mu}D^{ab}_{\mu}c^{b}\bigg)    \nonumber
\] 
and where we have introduced the operator $(A^h_\mu A^h_\mu)$ through the mass parameter $m^2$. Also, the transversality of $A_\mu^h$ is enforced by the 
Lagrange multiplier $\tau^{a}$. \\\\The action (\ref{sact}) is still non-local, since the expression for $(A^h_\mu A^h_\mu)$  is an infinite sum of nonlocal
terms.  Nevertheless, expression (\ref{sact}) can be cast in local form  \cite{Fiorentini:2016rwx}  by means of 
the introduction of an auxiliary localizing Stueckelberg field $\xi^a$, whose role is to give, for each gauge 
field $A_\mu$, its corresponding configuration that minimizes the functional $A^2$, {\it i.e.}, 
$A^h_\mu$. This is most naturally implemented by defining a field $h$
which effectively acts on $A_\mu$ as a gauge transformation would act, in order to provide the 
minimizing configuration $A^h$, that is,
\begin{equation}\label{local_Ah}
A^{h}_{\mu} \equiv A^{h,a}_{\mu}\,T^{a}=h^{\dagger}A_{\mu}h+\frac{i}{g}h^{\dagger}\partial_{\mu}h.
\end{equation}
with
\begin{equation}
h=e^{ig\xi}=e^{ig\xi^{a}T^{a}},     \label{hxi}
\end{equation}
where $\{T^a\}$ are the generators of the gauge group $SU(N)$ and $\xi^{a}$ is a Stueckelberg field. 
\\\\Thus, by substituting the expression (\ref{local_Ah}) for $A^h$ in the action (\ref{sact}), 
we now have a local theory in terms of the field $\xi$. The price 
one has to pay to have such a local theory is a non-polynomial action. Indeed, by expanding 
(\ref{local_Ah}), one finds an infinite series whose first terms are
\begin{equation}\label{Ah_expansion}
 (A^{h})^{a}_{\mu}=A^{a}_{\mu}-D^{ab}_{\mu}\xi^{b}-\frac{g}{2}f^{abc}\xi^{b}D^{cd}_{\mu}\xi^{d}
 +\mathcal{O}(\xi^{3})\,,   
 \end{equation}
where 
\begin{equation}
 D_\mu^{ab} = \delta^{ab}\partial_\mu - gf^{abc}A_\mu^c 
\end{equation}
is the covariant derivative in the adjoint representation. 
\\\\The nonlocal expression (\ref{ah}) for 
$A^h_\mu$ in terms of the gauge field $A_\mu$  can 
be recovered by imposing the transversality condition $\partial_\mu A^h_\mu=0$, {\it i.e.} after taking 
the divergence of both sides of (\ref{Ah_expansion}), equating it to zero and solving for the Stueckelberg 
field $\xi^a$ \cite{Capri:2015ixa,Capri:2015nzw,Fiorentini:2016rwx}. Due to the transversality condition enforced by the Lagrange multiplier $\tau^a$,  the Stueckelberg field 
$\xi^a$ acquires now a specific meaning:  it is precisely the field which brings a generic gauge 
configuration $A_\mu$ into the gauge-invariant and transverse field configuration $A^h_\mu$ 
which minimizes the functional $A^2_{min}$. As shown in \cite{Fiorentini:2016rwx},  this feature,  
encoded in the term $\int d^4x\;\tau^{a}\,\partial_{\mu}A^{h,a}_{\mu}$, gives rise to deep differences 
between our construction and the standard Stueckelberg mass term, which is known to be a 
non-renormalizable theory which has to be treated as an effective field theory  \cite{Ferrari:2004pd}. 
\\\\An important property  
of $A^h_\mu$, as defined by eq.(\ref{local_Ah}), is its gauge invariance, that is,
\begin{equation}
A^{h}_{\mu} \rightarrow A^{h}_{\mu} \;,
\end{equation}
as can be seen from the gauge transformations  with $SU(N)$ matrix $V$
\[
A_\mu \rightarrow V^{\dagger} A_\mu V + \frac{i}{g} V^{\dagger} \partial_ \mu V
\;, \qquad h \rightarrow V^{\dagger} h    
\]
The local version of the action (\ref{sact}), in terms of the Stueckelberg field 
$\xi^a$, is thus given by
\begin{eqnarray}
S &=& S_{FP}+
\int d^4x \left(\tau^{a}\,\partial_{\mu}A^{h,a}_{\mu}
+\frac{m^{2}}{2}\,A^{h,a}_{\mu}A^{h,a}_{\mu}\right)\nonumber\\
&=& S_{FP}+\int d^4x \left[ \tau^a\left(A^{a}_{\mu}-D^{ab}_{\mu}\xi^{b}\right)\right]
 \nonumber\\
&&+ \;\;\frac{m^2}{2}\int d^4x \left(A^{a}_{\mu}-D^{ab}_{\mu}\xi^{b}\right)
\left(A^{a}_{\mu}-D^{ae}_{\mu}\xi^{e}\right)\nonumber\\
&& + \;\;\;\cdots
\label{S_local}
\end{eqnarray}
Due to the use of the auxiliary Stueckelberg field $\xi^a$, expression (\ref{S_local}) exhibits a non-polynomial character. 

\subsection{BRST invariance}

The local action $S$ enjoys an exact BRST symmetry \cite{Fiorentini:2016rwx}:
\begin{equation}
s S = 0 \;, \label{brst_s}
\end{equation}
where the nilpotent BRST transformations are given by 
\begin{eqnarray}
sA^{a}_{\mu}&=&-D^{ab}_{\mu}c^{b}\,,\nonumber \\
sc^{a}&=&\frac{g}{2}f^{abc}c^{b}c^{c}\,, \nonumber \\
s\bar{c}^{a}&=&ib^{a}\,,\nonumber \\
sb^{a}&=&0\,, \nonumber \\
s \tau^a & = & 0\,, \nonumber \\
s^2 &=0 &\;.    \label{brst}
\end{eqnarray}
From \cite{Dragon:1996tk}, for the Stueckelberg field we have,  with $i,j$ indices associated with a generic representation,
\begin{equation}
s h^{ij} = -ig c^a (T^a)^{ik} h^{kj}  \;, \qquad s (A^h)^a_\mu = 0  \;,  \label{brstst}
\end{equation}
from which the BRST transformation of the field $\xi^a$ can be evaluated iteratively, yielding
\[
s \xi^a=  - c^a + \frac{g}{2} f^{abc}c^b \xi^c - \frac{g^2}{12} f^{amr} f^{mpq} c^p \xi^q \xi^r + O(\xi^3)  
\] 
As shown in \cite{Fiorentini:2016rwx}, the BRST invariance of the action $S$, eq.(\ref{sact}),  can be translated into 
functional identities which can be used to show that $S$ is in fact renornalizable to all orders of 
perturbation theory.

\section{Conclusion} 

A gauge invariant dimension two operator can be introduced by minimizing  the functional
$\mathrm{Tr}\int d^{4}x\,A_{\mu }^{u}A_{\mu }^{u}$ along the gauge orbit, {\it i.e.}
\begin{equation}
A_{\min }^{2} =\frac{1}{2} \mathrm{Tr}\int d^{4}x\,A_{\mu }^{h}A_{\mu }^{h}\;,   \label{cc1}
\end{equation}
with $A_\mu^h$ the transverse configuration, $\partial_\mu A^h_\mu=0$, given in expression (\ref{ah}). \\\\Despite the highly non-local character, a fully local set up for both operators 
$(A^h_\mu A^h_\mu)$ and $A^h_\mu$ can be constructed, giving rise to a local and BRST-invariant action $S$, eq.(\ref{sact}), which turns out to be renormalizable to all orders of perturbation theory \cite{Fiorentini:2016rwx}.  Let us conclude by mentioning that, owing to the gauge invariance of $(A^h_\mu A^h_\mu)$ and $A^h_\mu$, the corresponding anomalous dimensions, $(\gamma_{(A^{h})^2}, \gamma_{A^{h}})$, turn out to be independent from the gauge parameter $\alpha$ entering the gauge fixing condition, 
 being given by \cite{Fiorentini:2016rwx}
\begin{eqnarray}
\gamma_{(A^{h})^2} &= &  \gamma_{A^{2}}\Big|_{\rm Landau}  =-\left(\frac{\beta(a)}{a}+\gamma^{\mathrm{Landau}}_{A}(a)\right)    \nonumber \\
\gamma_{A^{h}}& = & \gamma_{A^{h}}\Big|_{\alpha=0}=\gamma^{\mathrm{Landau}}_{A}(a)  \nonumber \\
 a&=& \frac{g^{2}}{16\pi^{2}}\ \label{ad} 
\end{eqnarray} 
where $(\beta(a), \gamma^{\rm Landau}_A(a))$ denote, respectively, the $\beta$-function and 
the anomalous dimension of the gauge field $A_\mu$ in the Landau gauge, corresponding to set the gauge parameter $\alpha$ to zero, $\alpha=0$.    One sees therefore that $(\gamma_{(A^{h})^2}, \gamma_{A^{h}})$ are not independent parameters of the theory.

\end{document}